OpenStudy (acacia):

What is the equation of the ellipse with vertices at (-25, 0), (25, 0) and co-vertices (0, -15), (0, 15)

OpenStudy (acacia):

\[a. \frac{ x^2 }{ 625 }-\frac{ y^2 }{ 225 }=1\] \[b. \frac{ x^2 }{ 225 }+\frac{ y^2 }{ 625 }=1\] \[c. \frac{ x^2 }{ 225 }-\frac{ y^2 }{ 625 }=1\] \[d. \frac{ x^2 }{ 625 }+\frac{ y^2 }{ 225}\]